TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-03-19 23:24:00
subject: 3\07 Pt-1 ISS On Orbit Status 07-03-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
3\07 ISS On-Orbit Status 07-03-2003
Part 1 of 3

ISS On-Orbit Status 3/7/03

All ISS systems continue to function nominally, except as noted 
previously or below.

FE-1 Nikolai Budarin, assisted by CDR Ken Bowersox, uninstalled and 
removed the Kurs-A rendezvous and approach radar system from the 
10P/Progress M-247 motion control and navigation system (SUDN), a 
three-hour job. [These valuable components will be returned to Earth 
for "recycling", but only after Shuttle flights have resumed. Until 
then, the possibility of procuring a new Kurs system from the 
Ukrainian manufacturer is one of the issues our partner Rosaviakosmos 
(RSA) is faced with at this time.]

For FE-2/SO Don Pettit, troubleshooting of the MSG (Microgravity 
Science Glovebox) continued. [Today, Pettit re-installed another 
connector (P121) for the ground to monitor ESEM3 (exchangeable 
standard electronic module 3) current draw with power applied to the 
Left Filter Tube Sensor, the Left Delta P Sensor Box and the MTL 
(moderate temperature loop) IN-Temperature Sensor Conditioner. 
Afterwards, Don was to rotate the rack again and connect the next 
cable (P106) for another round of troubleshooting. POC (Payload 
Operations Center) is still evaluating the data gathered so far and 
hopes to have an MSG update soon.]

The crew completed a scheduled review of the uplinked Stage EVA2 
timeline and activities overview. Later in the day, they tagged up 
with EVA specialists in Houston to discuss the upcoming spacewalk 
tasks. A "refresher course" on fluid QDs (quick disconnects) and their 
corrective SPDs (spool positioning devices) is scheduled for Monday 
(3/10). Final go-ahead for the EVA is expected by the MMT (Mission 
Management Team) on 3/11, followed by formal EVA readiness review on 
3/18. [The EVA will probably be no earlier then 3/25; however, more 
detailed planning is required since the solar Beta angle will peak at 
that time (52 deg), necessitating some "interesting" tradeoffs between 
environmental parameters with EVA impacts. Objectives of the EVA are 
(1) S0/S1/P1 truss segment tasks like reconfiguring electrical 
connectors to provide additional inhibit to prevent SSAS 
(segment-to-segment attachment system) release, and NTA (nitrogen tank 
assembly) inspection to troubleshoot its heater problem; (2) R&R of 
the failed MT RPCM 3A (mobile transporter/remote power controller 
module #3A); (3) reconfiguring of CMG (control moment gyroscope) 
connections to ensure that no single EPS (electrical power system) 
failure will result in the loss of two of the three operating CMGs; 
(4) installation of two SPDs on Lab HX (heat exchanger) fluid QDs; (5) 
re-installing and securing of a QD thermal cover on the S1 RBVM 
(radiator beam valve module); (6) freeing the S1 CETA (crew equipment 
translation aid) railcart light stanchion and completing the light 
installation aborted during Stage EVA1; (7) and get-ahead tasks as 
time permits, such as tool retrievals and photo ops.]

Budarin started another round of routine 24-hr. checkup and watering 
of the Russian BIO-5 Rasteniya-2/Lada-2 ("Plants-2") plant growth 
experiment which he had set up yesterday. [The regular daily 
maintenance of the powered payload consists of monitoring seedling 
growth, humidity measurements, moistening of the substrate when 
required, and photo/video recording.]

Don Pettit completed the daily routine maintenance of the SM's SOSh 
life support system, including ASU toilet facilities, and later 
prepared the IMS inventory delta file for downlink. Bowersox conducted 
the regular status checkup of the autonomous PCG-STES010 payload in 
the Lab.

Nikolai Budarin meanwhile conducted the weekly inspection of the BRPK
air/condensate separator of the SRVK water processing system.

In the continuing assessment of Station resources, Nikolai later in 
the day did an inventory audit of Russian DVCAM and LIV audio/video 
equipment such as cables, MiniDV tapes and plug-in adapters. More 
audits to come. [Some of the items, which he was to search for, are 
listed in the IMS database as "lost".]

Bowersox prepared the data from the 5th Foot (foot/ground reaction 
forces during space flight) session yesterday (3/6) and transferred 
the file for subsequent downlink to MCC-H.

Early in the morning (2:37am EST), amateur radio operator Don Pettit 
chatted with students of Istituto Tecnico Industriale (ISIS) Malignani 
in Cervignano del Friuli, Northern Italy, a technical high school with 
more than 300 students directly involved with the amateur radio 
contact.

Later (11:58am), Pettit supported a second ham radio pass, this time 
with students of Krueger School of Applied Technologies (K-SAT) in San 
Antonio, TX, with a curriculum that features hands-on learning 
centered on aerospace.

At 9:36am, Bowersox and Pettit participated in two live TV-interactive 
PAO interviews of about 10 minutes each, with KPAM Radio, Portland, 
OR, and with the Indianapolis Star.

All crewmembers performed their regular daily physical exercise on 
TVIS (treadmill with vibration isolation and stabilization), RED 
(resistive exercise device), CEVIS (cycle ergometer with vibration 
isolation) and VELO bike with load trainer.

In a crew telecon, exercise specialists discussed possible approaches 
to troubleshoot the CEVIS, which has its control panel failed and can 
only be used in the unpowered mode. [No spare parts are being 
manifested on the upcoming Progress resupply flight, but it is hoped 
that troubleshooting by the crew can solve the problem.]

For the RED, Bowersox has advised against a suggested configuration 
change involving installation of the used Flexpack canister #1010, due 
to potential asymmetric loading. RED engineers uplinked a number of 
questions to better characterize the current status of this canister 
for closer analysis.

In the Russian segment (RS), both SM computers (TVM/Terminal and
TsVM/Central) continue on two out of three mutually redundant 
channels. Full reinitialization/reconfiguration is scheduled for 3/17, 
right before they receive the 7.01 software upload which is currently 
on hold. This date depends on the restoration of GNC (guidance, 
navigation & control) functionality in the U.S. segment (USOS) with 
the second R3 software patch.

The new GNC patch will be delivered to MCC-H by tonight. The mandatory
mission configuration (uplink) test is scheduled for Monday (3/10). If
successful, installation of the patch will follow on 3/12 in 
conjunction with the handover of ISS attitude control to the RS for 
the Progress thruster testing.

The test burns of the two Progress 10P thruster manifolds are set for 
3/12 (5:23pm EST) and 3/13 (6:38pm) next week, each one a mini-reboost 
of 1.38 m/sec delta-V for the first, 0.31 m/sec for the second burn.

The next larger Station reboost, of about 6 m/sec, is scheduled for 
4/14.  Its objective is to adjust the ISS in-orbit phasing angle for 
the launch of 6S/Soyuz TMA-2 in early May and the next two Soyuz 
flights as well (7S and 8S). [The 4/14 burn is part of the longer-term 
altitude strategy being developed by Moscow since after the Shuttle 
stand-down. The tentative strategy extends to December 2004 and is 
based on the "most pessimistic" assumptions, including ULF-1 not 
flying before August 2004. Under these conditions, best 
propellant-savings are obtained by early maximizing ISS orbit altitude 
to the extent allowed by Soyuz and Progress capabilities.]

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.